A 10 foot ladder is leaning against a vertical wall. The base of the ladder is 3 feet from the wall. How high is the point where the ladder touches the wall?
Answer by nerdybill(7384)

Apply Pythagorean theorem:
Length of the ladder is the hypotenuse.
.
Let x = height where the ladder touches the wall
3^2 + x^2 = 10^2
9 + x^2 = 100
x^2 = 91
x = 9.54 feet